SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jan 17th, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Arik Kreisel, Selina Li, Rahmat Rahmat, Su Dong, Mark Tibbetts.

Operations

  • New trigger on-call cellphone expected this week. The appropriate databases will be updated upon it's arrival.
  • The old TSF channels will not be deleted from the EPICS database, they'll just stop being archived. They should still be deleted from the IOC database though.
  • The EPICS display seem to be showing nothing from the TSF, a problem which is unrelated to the changes made in the EPICS display. Where did the information go? Speak to Ray.
    • I (Debbie) mentionned this in the 3.45 meeting and Mike Przybylski talked to me about it - turns out when the CAN-BUS cables were reconnected after the old TSF crates were removed things went wrong. There was no terminator on the final TSF crate of the chain, which oddly wasn't a problem before but caused everything to break once the in/out oldTSF cables were connected. The new TSF was left out of the fixed chain, which is why nothing was seen in EPICS. It's now been added in again, see here to see how the optical power is now what we'd expect.

EMT

  • AOBs are being changed in the EMC tomorrow during the ROD, so Mark will keep an eye on that.
  • The turn-on threshold and slope will be studied, first by reproducing Dan's mumu plots with raw EMC energy, rather than converted ADC counts. Threshold and slope can then be compared MC to data, and raw energy compared to ADC count to confirm the conversion factor.

DCT/DCZ

  • Fang uploaded a new version of the ROM libraries to keep in line with the online relase 22.0.0 (16.0.0).
  • Once the DCH t-to-d constants have been uploaded, a DCH-DCT calibration should be taken. Selina showed Debbie how to do this.

IFT

  • The IFT time latency is now aligned for RPC and LST, as well as with other subsystems. See HN posted.

GLT

  • How to fix this bug that requires the ROM be rebooted if the crate is powercyled? There's a tag for the correct firmware somewhere, it should be looked up!

L3

  • Online release 22.0.0 (16.0.0) includes the new L3 code. Validation is underway to compare it to 18.7.1 - with 100,000 events there appeared to be some differences (<1%) with the hadronic and lumi filters, Rainer is running over 1e6 events to check. Any change in the Lumi filter is worrying! This release does not include the new configurable c-track table - it's the biggest change to be made but hasn't been signed off yet by the tracking group.
  • A first look at the t-to-d calibrations show constants that are very similar to the old ones - the only change seems to be in hit residuals which have become more gaussian in distribution. We should get more TSF W=3 segments with higher voltage and we need to upload the new constants to check all this, but there should be no change in L3 efficiency

Data Quality and Monitoring

  • The changes in the stripchart have been given the green light by the DQM group and will be implemented as soon as a few last bugs have been ironed out. Rahmat will post a link to the Run5b charts he showed before christmas.

Trickle

  • There was a meeting at 3pm to discuss the technical details of the reverse window trickle monitoring. A quantum change is understood - the deadtime scales with quanta, but and change in frequency would be more of a qualitative problem.

Config

  • Joe ran Moose with the tight/emergency configurations and ran into a few problems - the L3 trigger lines need ot be configured as well as L1 lines! This has been fixed, but the FCT prescale simulation does not support missing lines, which is a problem for the emergency configuration with it's deleted lines. So, for now prescale the "deleted" lines as much as possible, but leave them where they are to make the timing compatible.
  • Debbie made some plots to look at the 1Zk line, mainly to check whether the asymmetric momentum cut was applied in the correct charge direction. The plots are here. The show Pt for single track MC electrons ("en") and positrons ("ep") in the momentum range 0.5-3.5GeV, and for single track K+ ("kp") and K- ("kn") in the momentum range 0.1-1.1GeV. The red line is Pt (taken from DCH information and multiplied by charge) for all tracks, and the black line is those tracks for which 1Zk fired. It appears that there is a larger affect on the positive particles than the negative particles, which is what is expected if the asymmetric cut is correctly applied. However, muons are the best things to look at, the truth information is needed and the MC generated spectrum of momentum should be looked at as well - is it flat? Coming soon...

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.